Journalist Report – February 27th

Crew 206 – 02/27/2019

Sol 3

Author : Auzou Benjamin, Journalist

"Martian Excursion"

We are born to explore; since the dawn of humanity, humans never stopped
to explore new places. The first humans settled to all five continents,
the navigators of the XVth century sailed all seas, the astronauts of
the XXth century flew to space. Today was an exploration day on Mars !
The team of the EVA (Aurélien, Norbert, Gaspard and I) rode the rovers
to the Moons area. We walked about an hour on the muddy and slippery
slopes of the White and Grey Moons, following the steps of Norbert, our
scout of the day, under the wise supervision of the EVA leader Aurélien.

To avoid fatigue after lunch, I visited Norbert in the Green-Hab and
helped him taking care of radishes and carrots while some others were
taking a nap. Back in the Hab, Aurélien was waiting for us all to show
us his first night observation : Orion’s Nebula.

This afternoon we tested our own analog spacesuit in the MDRS for the
first time. This project was born in September 2017, when we first heard
about the MDRS and decided to apply to a rotation. We worked since that
moment on the conception and fabrication of this spacesuit. That’s why
we are happy to see this project concretise here. The prototype is far
from what we expected but we are focused on our goal to take some of us
in EVA with this suit and we will make all changes and adjustments
needed to reach it.

Today we had the first results of our strict water monitoring. And this
is a huge improvement compared to the monitoring of crew 189, the last
mission of ISAE-SUPAERO students. For Sol 2, the total consumption is
77.4 liters. The Green-Hab is the first source of consumption with 41%,
then come drinking water with 25%, the flushes with 13%, cooking water
and dishes water with 5% and 6%. The last 8% are unmeasured and come
from leaks or flushes volume uncertainties. In comparison, the average
unmeasured water for crew 189 was 37% of total consumption.

For more than a year, we worked together on the preparation of this
mission. We passed through happy and bad moments, funny talks and long
crisis meetings. But now we are here, on Mars, and we are like a family,
forged by our common passion for space exploration, our interventions
with young students and this mission. What we are achieving now is
exciting and fulfilling !

Sol Summary – February 27th

Crew 206 February 27th 2019
Sol 3

Summary Title : Taking care of the Hab

Author : Jeremy Auclair

Mission Status : All nominal

Sol Activity Summary : Once again we got out of bed at 07:00 am to
workout, but everybody had slept well, we went to bed early yesterday.
Our Dreem Headbands showed that we had an effective time of sleep
ranging from 6h20 and 8h20, I slept more than everyone else ! After
breakfast we prepared the EVA. It was the first EVA without a veteran,
Aurélien was EVA leader. It went very well, Opportunity’s battery
decreased quickly so they went for plan B. They explored the moon areas.
During the EVA Cerise and I cleaned the Hab’s lower and upper deck. I
then took care of the water consumption data : on Sol 2 (yesterday) we
used 77,4L (20,5 Gallons), with 41% for the GreenHab, 25% for drinking,
13% for flushes, 6% for dishes, 5% for cooking, 2% for hygiene (no
showers yet) and 8% of unmeasured consumption. Unmeasured consumption
comes from inaccuracies and small mistakes in counting, this part of
daily consumption should decrease during the mission. After lunch,
everyone worked on their experiments (except me, I took a good nap).
Aurélien, Benjamin, Gaspard and Norbert did some pre-EVA tests on their
space suit, all the vital functions work, but it still needs to be
improved the ergonomic standpoint. Cerise worked on Aquapad to measure
the microbial quality of our water. First results suggest that the water
drinkability is problematic, but it is surely a manipulation or data
interpretation, we have sent a mail to the BioMérieux laboratories to
have an explanation. Before the COMMS, as planned, we participated to the
Virtual Reality experiment.

Look ahead plan : Tomorrow is once again going to be a busy day, we have
4 experiments planned. Tomorrow’s EVA will allow us to train the last
members of the crew.

Anomalies in work : None

Weather : Mostly cloudy

Crew physical status : Fine

EVA : Beige moon area

Reports to be filed : GreenHab, Journalist, Operation, EVA Report, EVA
Request, Astronomy

Support Requested : None

Operations Report – February 27th

Crew 206 Operations Report 27 – FEB – 2019

SOL: 3

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need
replacement on generator. A contractor has been arranged to make these
repairs within a few days.

Generator (hours run): Turned off at 0800 (27th Feb), Turned on at 1840
(27th Feb)

Solar— SOC % – Turned off Gen (24th Feb) 64%, Turned on Gen (24th Feb) 59%

Diesel Reading – 3/4 (on a scale of 0 – 1/4 – 1/2 – 3/4 – 1)

Propane Reading – 21 (± 2)

Ethanol Free Gasoline – 7 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – around 500 gallons (90 %)

Water differential (Static tank) – 12 gallons (± 2 gallons)

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – 276.5 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01415960 units

Toilet tank emptied: Yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Used

Hours: 52.4 hours

Beginning charge: 100%

Ending charge: 34%

Currently charging: Yes

Curiosity rover used: Used

Hours: 76.7 hours

Beginning charge: 100%

Ending charge: 61%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 78.3 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Opportunity’s battery decreased faster
under 60%, and lasted way less than Curiosity’s. (for more
detailed battery readings during the EVA, cf EVA report)

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Used for Staff

Staff Crew Car used and why, where? Off Campus

General notes and comments:

Summary of internet: Lost Internet access a lot again this
morning. Did a software reset on the modem. Working
without interruption now.

Summary of suits and radios: Suit #9 Fixed, currently charging. Suit #2
experienced fogging. Checked it, seems ok

Summary of Hab operations: Hotfixed the lamp bulb in upper
deck. It is stable now. Inspection revealed the plastic bulb is broken on
several loactions.

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAM operations:

Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing more to report.

EVA Report – February 27th

Crew 206 EVA Report 27-Feb-2019
EVA #3

Author: Aurélien Mure

Purpose of EVA: The EVA is built to begin the exploration and to
familiarize the crew with mid range EVA’s. We will go with the rovers to
the moons region and start hiking from this point. The purpose will be
to explore by walking The Moon, Gray Moon, Beige Moon and Yellow Moon.
The rovers will be parked at the intersection between Copernicus Hwy and
Brahe Hwy, the rest will be done by walking counterclockwise starting
with Yellow Moon. It is important for us to start doing long EVA’s to be
used to it quickly. As 2 of us will be commanders of crews in 2020 and
we can’t know yet who would it be, the whole group needs to be trained
to all types of EVA’s. That is why we will continue to ask for EVA’s
with four marsonauts. Moreover, in our mind being two on each rover is
safer than being alone.

If we realize that the batteries won’t allow this, we will stop at White
moon and explore by hiking Beige Moon and Gray Moon, or if we have to
stop before near Reservoir Dam and Cowboy Corner.

Start time: 9:08 am

End time: 11:14 am

Narrative: After the 5 minute depressurization, our team Blue (Gaspard
and me) checked the rovers state, they were charged and ready. Team Red
(Benjamin and Norbert) checked the water level of the static tank. Team
Blue will ride Opportunity and Team Red Curiosity. The departure was at
9:18. We drove to the LOAC place to check quickly the battery level.
12,2V, everything is nominal, our experiment is on its ay. We drove
carefully with as objective saving energy. We checked the battery every
2 min to decide if we went to our plan A (Copernicus Hwy), plan B
(beginning of Brahe Hwy) or plan C (Reservoir Dam). We went through our
plan C destination with enough energy to continue until plan B
destination. Just after Reservoir Dam we decided to focus on plan B. We
arrived at the beginning of Brahe Hwy with 67% of battery for
Opportunity and 72% for Curiosity. We went for a walk in the region of
Grey Moon, for an hour and around 3km. The ground was very slippery so we
were extremely careful with our steps.

At 10:46 we decided to leave the place and go back to the Hab. We
arrived at 11:06, with Curiosity at 61% and Opportunity at 34%. With
Spirit that is supposed to be the laziest one, we stick to the reality
as on mars Curiosity is in operation, Opportunity has died 4 months ago
and Spirit a long time ago !

Destination: Gray Moon
Coordinates : 516450/4254800

Participants: Aurélien Mure (XO, EVA leader), Benjamin Auzou (COMMS ),
Gaspard Thieulin (ENG), Norbert Pouzin (GHO)
Road(s) and routes per MDRS Map : Entrance Road , Cow Dung Road

Mode of travel: Driving and Walking

Greenhab Report – February 27th

Crew 206 Green Hab Report – 27-FEB-2019

GreenHab Officer: Norbert Pouzin

Environmental control: Heating and Cooling w/ ambient air (0 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 18.2°C

High: 35.9°C

Hours of supplemental light: 7 Hours

Daily water usage for crops: 8.6 gallons

Water in Blue Tank: 276.5 gallons

Crops watered at 0940 hours (2.2 gallons) and 1830 hours (6.4 gallons):

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)

1) 09.40 AM 35.0° C / 0 %
2) 11.30 AM 24.8° C / 0 %
3) 15:00 PM 21.2° C / 24 %
4) 18:30 PM 22.8° C / 33 %

Changes to crops: I replanted radishes.
Radishes from my experiment are starting sproutling.
Carrots replanted yesterday don’t all look great.
Tomatoes look great, but in two pots they haven’t germinated.

Narrative:
This morning when Jeremy took care of the greenhab (I was in EVA), it
was really hot. We had difficulties to monitor the heat of the hab in
the morning : the fan cooled too much the greenhab.
The door was opened from 0940 to 1630 ; the fan from 1100 to 1130.
I replanted some radishes.

Harvest: None.

Support/supplies needed: None.

Astronomy Report – February 27th

Astronomy Report, Sol 3
Name: Aurélien Mure
Crew: 206
Date: 02/25/2019
ROBOTIC OBSERVATORY
Object observed : Orion Nebula, M42. I attach the image I did from the observation. I’m a beginner in astrophotography, I hope my pictures will be improved during the mission.
Problems encountered : I think I exposed to much the telecope, the center of the nebula is to bright.